HIGH-RESOLUTION INFRARED SPECTRA OF JET-COOLED METHOXY RADICAL (CH_{3}O): C-H STRETCH VIBRATIONS

\underline{JIA-XIANG HAN}, Y. G. UTKIN, HONG-BING CHEN, L. A. BURNS, AND R. F. CURL, Rice University, Department of Chemistry and Rice Quantum Institute, Houston, Texas 77005.

The high-resolution infrared spectra of jet-cooled methoxy radical (CH_{3}O) in the region 2850-2940 cm^{−1} have been recorded in pulsed slit jet expansions with the excimer laser flash photolysis method. A perpendicular band centered at 2885.7 cm^{−1} was assigned to a portion of the C-H asymmetric stretch vibration of CH_{3}O using ground state combination differences. Five sub-bands with P′′ = -3/2, -1/2, +1/2, +3/2, +5/2 and J′′ up to 8.5 have been assigned. The series of Q branches almost equally separated by 12.4 cm^{−1} with intensity patterns characteristic of perpendicular transitions supports the assignment. Because of the low jet temperature, the assigned transitions all arise from the E_{3/2} fine structure component of the ground state. The asymmetric C-H stretch vibration is subject to a strong Jahn-Teller effect, which should produce several subbands. The firm vibrational assignment of this band requires data for all C-H stretch bands. Least squares fitting of this band and scanning over more of the C-H stretching region scan for other bands are proceeding in our lab.